# Strongly magnetized accretion with low angular momentum produces a weak jet

### v1

@@ Abstract
We study spherical accretion of magnetized plasma with low angular momentum onto a supermassive black hole, utilizing global General Relativistic Magnetohydrodynamic simulations. Black hole-driven feedback in the form of magnetic eruptions and jets triggers magnetized turbulence in the surrounding medium. We find that when the Bondi radius exceeds a certain value relative to the black hole’s gravitational radius, this turbulence restricts the subsequent inflow of magnetic flux, strongly suppressing the strength of the jet. Consequently, magnetically arrested disks and powerful jets are not a generic outcome of accretion of magnetized plasma, even if there is an abundance of magnetic flux available in the system. However, if there is significant angular momentum in the inflowing gas, the eruption-driven turbulence is suppressed (sheared out), allowing for the presence of a powerful jet. Both the initially rotating and non-rotating flows go through periods of low and high gas angular momentum, showing that the angular momentum content of the inflowing gas is not just a feature of the ambient medium, but is strongly modified by the eruption and jet-driven black hole feedback. In the lower angular momentum states, our results predict that there should be dynamically strong magnetic fields on horizon scales, but no powerful jet; this state may be consistent with Sgr A* in the Galactic Center.
@@ 1 Introduction
General relativistic magnetohydrodynamic (GRMHD) simulations are frequently used to study accretion flows around supermassive black holes (BHs) (e.g., Porth et al., 2019). In the context of low luminosity accretion flows, radiation is dynamically unimportant, which leaves ideal GRMHD method scale-free. Typically, these models are initialized with a rotating Fishbone-Moncrief torus (Fishbone & Moncrief, 1976), which is threaded with a dynamically weak magnetic field organized on large scales. In the absence of magnetic fields, this represents an equilibrium solution.
The dynamics of magnetized accretion flows typically fall into two categories: Magnetically Arrested Disk (MAD) and Standard and Normal Evolution (SANE) (Narayan et al., 2012). The MAD state is characterized by strong coherent magnetic fields that become dynamically important, effectively regulating the accretion flow by creating magnetic forces strong enough to counterbalance the gravitational pull of the BH. Specifically, the magnetic flux threading the BH’s horizon, $\Phi_{\rm BH}$, is argued to saturate at a maximum value for a given mass accretion rate, $\dot{M}_{\rm BH}$. The MAD flux is found to be approximately $50\sqrt{\dot{M}_{\rm BH}r_{g}^{2}c}$, where $r_{g}$ is the gravitational radius of the BH and $c$ is the speed of light. When the magnetic flux at the horizon exceeds the MAD limit, some of the flux erupts and escapes from the BH through magnetic reconnection (Ripperda et al., 2022), leading to episodic fluctuations in $\Phi_{\rm BH}$. Most importantly, the MAD state is associated with the production of relativistic jets with electromagnetic luminosity exceeding the accretion power, $\dot{M}_{\rm BH}c^{2}$ (Tchekhovskoy et al., 2011).
@@
Ressler et al. (2021) showed that a MAD could also form as a result of a low-angular momentum magnetized flow onto a spinning BH – an idealization of the stellar wind accretion problem. They examined various orientations of the initial magnetic field and found that the dynamics of the flow near the horizon closely resemble MAD dynamics. However, on larger scales, they argued that the jet is unstable due to the current-driven kink instability. A similar problem was investigated by Lalakos et al. (2024). By increasing the Bondi radius by a factor of 10 compared to Ressler et al. (2021), and focusing on a magnetic field aligned with the BH’s spin, they also found that the system quickly transitions to a MAD state. However, as time progressed, their simulation transitioned into a state without a powerful jet, which they argued to be a result of the jet becoming kink-unstable (by criteria described in Tchekhovskoy et al. 2011; Bromberg & Tchekhovskoy 2016) and disrupting the accretion flow. Thus, two different regimes, resulting from different Bondi radii, were identified.
@@
Studying a similar problem, Kwan et al. (2023) conducted a series of GRMHD simulations of spherical accretion of magnetized medium with varying initial angular momentum. In contrast to Ressler et al. (2021) and Lalakos et al. (2024), they argued that no jet can form in the zero angular momentum case. The inclusion of angular momentum resulted in a powerful jet capable of propagating to large distances. They attributed this to the dependence of the efficiency of the angular momentum transport on the magnetorotational instability.
@@
The diversity and seeming incompatibility of the results in the literature on the properties of jets in low angular momentum flows motivates us to reexamine this problem. Spherical accretion of magnetized plasma is also the simplest problem in which to understand magnetic flux accumulation by a central black hole, a key problem in accretion theory given the central role of magnetic fields in angular momentum transport and jet/outflow production. The most idealized version of this scenario involves only a few key parameters: the spin of the BH, the Bondi radius of the surrounding medium, the strength of the magnetic field at the Bondi radius (characterized by plasma-$\beta$) and its orientation relative to the BH spin, and the initial angular momentum (or, equivalently, circularization radius) and its direction. In this study, we restrict our analysis to the case of aligned magnetic, rotation, and BH spin directions but study the effects of different Bondi radii, initial magnetic field strengths, black hole spin, and initial gas angular momentum.
@@
The remainder of the paper is organized as follows. We describe the setup of the GRMHD simulations used in this work in § 2 and present our results in § 3. We first provide an illustrative example of the dependence of the magnetized spherical accretion problem on the chosen Bondi radius in § 3.1. Then, we discuss the dependence of our results on the key parameters of our study in § 3.2; we consider the case of finite initial angular momentum – and the time-evolution of the resulting angular momentum – in § 3.3.
@@ 2 Method
We conduct GRMHD simulations utilizing the GPU-accelerated code H-AMR (Liska et al., 2019), which solves the MHD equations in a conservative form in curved space-time. Employing horizon-penetrating Kerr-Schild coordinates $r_{KS}$, $\theta_{KS}$, $\phi_{KS}$, we concentrate on a non-spinning BH characterized by a dimensionless spin parameter $a=0$, as well as a highly spinning BH with $a=0.95$. In the case of a highly spinning BH, its spin axis aligns with the polar axis. Our results, however, are insensitive to the inclination of the metric relative to the polar singularity of the coordinate system [^1] We demonstrate the insensitivity to the orientation with respect to the pole by tilting the BH spin by $90^{\circ}$ in Appendix A..
We conduct simulations at three resolutions, which we label as small, medium, and large. These correspond to effective grid sizes, $N_{r}\times N_{\theta}\times N_{\phi}$, of $384\times 144\times 256$, $812\times 384\times 512$, and $1536\times 768\times 1024$ cells. We employ “derefinement” near the pole, which efficiently reduces computational costs while maintaining a comparable spatial resolution near both the pole and the equator. The outer boundary, $R_{\rm out}$, is set to $10^{4}r_{g}$ in all simulations unless stated otherwise (the outer boundary of one of the simulations was extended to $10^{5}r_{g}$ to accommodate a Bondi radius of $5000r_{g}$). Our primary focus will be on the medium-resolution runs, as we observe convergence of the main results with resolution (see Appendix A). Throughout the paper, we measure distance and time in $r_{g}$ and $r_{g}/c$ respectively.
@@
where $c_{s}$ represents the initial gas sound speed. Initially, a zero angular momentum fluid with $\gamma=5/3$ is distributed throughout the simulation domain, characterized by a uniform density $\rho_{0}$ (we set $\rho_{0}=1$ in code units). The internal energy density is also uniform and determined by $r_{b}$:
@@
We set $\beta_{0}=10^{2}$ in the initial conditions of our $r_{b}=100r_{g}$ simulation. In a magnetized Bondi solution with a radial monopole magnetic field, $B^{r}$ undergoes stretching, $B^{r}\propto r^{-2}$. Therefore, to maintain similarity in the plasma-$\beta$ and magnetic flux ($\Phi_{\rm BH}$, described below) near the BH across different runs, we scale it according to the Bondi profile, $\beta\propto r^{3/2}$. Thus, we set $\beta_{0}=10^{3}$ for $r_{b}=500r_{g}$ and $\beta_{0}=3\times 10^{3}$ for $r_{b}=1000r_{g}$. This choice is not critical, however, as our result are in fact insensitive to the specific $\beta_{0}$ value, as long as it is initially above unity. We validate this statement by considering 4 values of $\beta_{0}$ for $r_{b}=500r_{g}$ (see § 3.2): $10$, $10^{2}$, $10^{3}$, and $10^{4}$. This encapsulates our simulations’ three key parameters: spin $a$, Bondi radius $r_{b}$, and initial plasma-$\beta_{0}$. We also discuss the dependence of our results on the initial angular momentum of the gas, parameterized using the circularization radius, $r_{c}$, in § 3.3.
The Bondi accretion rate for an isothermal flow is:
@@ 3 Results
First, we focus on the zero initial angular momentum case and the illustrative comparison of $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$ (with $\beta_{0}=10^{2}$ and $\beta_{0}=10^{3}$ respectively) at $a=0$ and $a=0.95$ in § 3.1. These results are in fact more generic than it might seem because we find convergence of our results as we increase $r_{b}$ yet further and vary $\beta_{0}$ (§ 3.2). Finally, the impact of angular momentum is discussed in § 3.3.
@@ 3.1 Early evolution
In this Section, we discuss the time evolution of our four main cases. We then discuss the role of eruption episodes as well as the magnetic field evolution. The non-spinning case converges with time and reaches a steady state. The highly spinning case initially reaches a similar steady state. However, at a later time, the dynamics of the spinning case with a larger Bondi radius, $r_{b}=500r_{g}$, is governed by the angular momentum generated by the BH feedback; this will be discussed in more detail in § 3.3.
@@ 3.1.1 Temporal evolution
Initially, the accreting plasma undergoes free-fall onto the BH, bringing the magnetic flux towards the horizon. Consequently, with the increasing accretion rate at the horizon, plasma properties such as density and velocity conform to the expected Bondi profiles. As long as the magnetic field is dynamically subdominant, $B^{\theta}$ is advected by the radial velocity, while $B^{r}$ experiences stretching according to $B^{r}\propto r^{-2}$. Therefore, the total pressure (both thermal and magnetic components) for a Bondi-like flow profile is given by:
$P_{\rm tot}(r,\theta)=\frac{2\rho_{0}M}{\gamma r_{b}}\left[\left(\frac{r_{b}}{r}\right)^{5/2}+\frac{1}{\beta_{0}}\left[\sin^{2}\theta+\cos^{2}\theta\left(\frac{r_{b}}{r}\right)^{4}\right]\right].$ (9)
For $\beta_{0}\gg 1$, this yields $P_{B}\ll P_{\rm th}$ close to the equatorial plane $\theta=\pi/2$. At higher altitudes, close to the BH, $P_{B}$ exceeds $P_{\rm th}$ due to its stronger dependence on radius compared to $P_{\rm th}$.
Figure 1: Comparison of the time evolution of (a) the accretion rate $\dot{M}/\dot{M}_{B}$; (b) magnetic flux $\Phi_{\rm BH}$; (c) normalized dimensionless magnetic flux $\phi_{\rm BH}$, and (d) outflow energy efficiency $\eta$ for different $r_{b}$ ($100r_{g}$ and $500r_{g}$ are shown by different colors) and spin values ($a=0$ and $a=0.95$ are represented by thin and thick lines respectively). All quantities are measured at $5r_{g}$. On the right panel (e-h), a zoom-in of panels (a-d) is shown. Note different x-axes for $r_{b}=100r_{g}$ (bottom x-axis) and $r_{b}=500r_{g}$ (top x-axis), represented by the corresponding colors.
@@
At a smaller Bondi radius (green lines), the dimensionless magnetic flux reaches a value indicative of the MAD state, $\phi_{\rm BH}\approx 50$ (c). The energy outflow efficiency, $\eta$, is above $100\%$, demonstrating efficient jet production through the Blandford-Znajek mechanism for $a=0.95$ case (thick green line). For $r_{b}=100r_{g}$ and $a=0.95$, the oscillations in $\Phi_{\rm BH}$, $\phi_{\rm BH}$, and $\eta$ are similar to reconnection-powered eruptions found in MADs that are produced in simulations starting from rotating tori (f-h).
At a larger Bondi radius of $r_{b}=500r_{g}$, however, the quasi-state dimensionless magnetic flux is smaller than for $r_{b}=100r_{g}$. This trend persists for both non-spinning (thin lines) and spinning (thick lines) cases though the effect is stronger for the spinning BH (thick purple line saturates at a smaller value, compared to the thin purple line in panels b and c).
@@
At a later time, the spinning BH with a larger Bondi radius (thick purple line) shows an increased $\phi_{\rm BH}$ and efficient energy outflow $\eta$ in the form of a powerful jet. We find such periodic jets to be a consequence of a periodic increase of angular momentum in the inflow, which will be explored in more detail in § 3.3.
@@
Figure 2: Plasma-$\beta$ in the poloidal plane for runs with different initial parameters $r_{b}$ and $a$. The first two panels show $a=0$, and the last two show $a=0.95$. (a): $r_{b}=100r_{g},\ \beta=10^{2},\ a=0$; (b): $r_{b}=500r_{g},\ \beta=10^{3},\ a=0$; (c): $r_{b}=100r_{g},\ \beta=10^{2},\ a=0.95$; (d): $r_{b}=500r_{g},\ \beta=10^{3},\ a=0.95$. The snapshots are taken at a time $40000r_{g}/c$ for $r_{b}=100r_{g}$ and $120000r_{g}/c$ for $r_{b}=500r_{g}$.
@@
In Figure 2, we present a side-by-side comparison of plasma-$\beta$ in the same four simulations in the poloidal $r-\theta$ plane with magnetic field lines shown by black lines. The first two columns illustrate the case of a non-spinning BH with $r_{b}=100r_{g}$ (a) and $r_{b}=500r_{g}$ (b). Panels (c) and (d) showcase a highly spinning BH, $a=0.95$, again with $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$, respectively. In both cases where $r_{b}$ is smaller ($100r_{g}$, corresponding to higher gas temperature), a more laminar flow is observed compared to the scenario with $r_{b}=500r_{g}$. Moreover, in the case of a spinning BH, smaller $r_{b}$ shows the presence of a jet (c), which is not seen in the $r_{b}=500r_{g}$ case (d). The snapshots are taken at the time of $40000r_{g}/c$ and $82000r_{g}/c$ for $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$, respectively.
@@
Figure 3: Three-dimensional rendering of the late-time evolution of plasma-$\beta$ and magnetic field lines, plotted on a scale of $250r_{g}$ for $r_{b}=100r_{g}$ (top row, a-b) and $r_{b}=500r_{g}$ (bottom, c-f). The origin of magnetic field lines is fixed at two regions: around the BH and at $\approx 100r_{g}$ from the BH. The corresponding movies can be found for $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$.
@@
In Figure 3, we show a three-dimensional rendering of plasma-$\beta$ (in color) and magnetic field lines for $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$ at $a=0.95$ at a large scale of $500r_{g}$. The magnetic field lines originate from two regions: on a sphere around the BH and approximately $100r_{g}$ from the BH. In all snapshots, the spin axis of the BH points upwards. The top row (a-b) shows two snapshots for the case of $r_{b}=100r_{g}$. Both demonstrate that the powerful jet extends to large distances and is likely kink-unstable beyond the Bondi radius. Magnetic flux tubes, which are remnants of previous eruption episodes we will discuss in detail in the next section, are present around the BH. Panels (c-f) show four snapshots for the case of $r_{b}=500r_{g}$. At times of $30\times 10^{3}r_{g}/c$ (c) and $100\times 10^{3}r_{g}/c$ (e), the jet is weak and is significantly tilted relative to the vertically oriented spin axis of the BH. Between those times, a jet with $\eta\approx 100\%$ reappeared at a time of $60\times 10^{3}r_{g}/c$ (d). At a time of $160\times 10^{3}$ (f), $\phi_{\rm BH}$ increases again (as seen in Fig. 1), leading to a strong jet with $\eta>100\%$.

### v2

@@ Abstract
We study the spherical accretion of magnetized plasma with low angular momentum onto a supermassive black hole, utilizing global general relativistic magnetohydrodynamic simulations. Black hole-driven feedback in the form of magnetic eruptions and jets triggers magnetized turbulence in the surrounding medium. We find that when the Bondi radius exceeds a certain value relative to the black hole’s gravitational radius, this turbulence restricts the subsequent inflow of magnetic flux, strongly suppressing the strength of the jet. Consequently, magnetically arrested disks and powerful jets are not a generic outcome of the accretion of magnetized plasma, even if there is an abundance of magnetic flux available in the system. However, if there is significant angular momentum in the inflowing gas, the eruption-driven turbulence is suppressed (sheared out), allowing for the presence of a powerful jet. Both the initially rotating and nonrotating flows go through periods of low and high gas angular momentum, showing that the angular momentum content of the inflowing gas is not just a feature of the ambient medium, but is strongly modified by the eruption and jet-driven black hole feedback. In the lower-angular-momentum states, our results predict that there should be dynamically strong magnetic fields on horizon scales, but no powerful jet; this state may be consistent with Sgr A* in the Galactic center.
@@ 1 Introduction
General relativistic magnetohydrodynamic (GRMHD) simulations are frequently used to study accretion flows around supermassive black holes (BHs) (e.g., Porth et al., 2019). In the context of low-luminosity accretion flows, radiation is dynamically unimportant, which leaves the ideal GRMHD method scale-free. Typically, these models are initialized with a rotating Fishbone-Moncrief torus (Fishbone & Moncrief, 1976), which is threaded with a dynamically weak magnetic field organized on large scales. In the absence of magnetic fields, this represents an equilibrium solution.
The dynamics of magnetized accretion flows typically fall into two categories: Magnetically Arrested Disk (MAD) and Standard and Normal Evolution (SANE) (Narayan et al., 2012). The MAD state is characterized by strong coherent magnetic fields that become dynamically important, effectively regulating the accretion flow by creating magnetic forces strong enough to counterbalance the gravitational pull of the BH. In particular, the magnetic flux threading the BH’s horizon, $\Phi_{\rm BH}$, is argued to saturate at a maximum value for a given mass accretion rate, $\dot{M}_{\rm BH}$. The MAD flux is found to be approximately $50\sqrt{\dot{M}_{\rm BH}r_{g}^{2}c}$, where $r_{g}$ is the gravitational radius of the BH and $c$ is the speed of light. When the magnetic flux at the horizon exceeds the MAD limit, some of the flux erupts and escapes from the BH through magnetic reconnection (Ripperda et al., 2022), leading to episodic fluctuations in $\Phi_{\rm BH}$. Most importantly, the MAD state is associated with the production of relativistic jets with electromagnetic luminosity exceeding the accretion power, $\dot{M}_{\rm BH}c^{2}$ (Tchekhovskoy et al., 2011).
@@
Ressler et al. (2021) showed that a MAD could also form as a result of a low-angular-momentum magnetized flow onto a spinning BH – an idealization of the stellar wind accretion problem. They examined various orientations of the initial magnetic field and found that the dynamics of the flow near the horizon closely resemble MAD dynamics. However, on larger scales, they argued that the jet is unstable due to the current-driven kink instability. A similar problem was investigated by Lalakos et al. (2024). By increasing the Bondi radius by a factor of 10 compared to Ressler et al. (2021), and focusing on a magnetic field aligned with the BH’s spin, they also found that the system quickly transitions to a MAD state. However, as time progressed, their simulation transitioned into a state without a powerful jet, which they argued to be a result of the jet becoming kink-unstable (by criteria described in Tchekhovskoy et al. 2011; Bromberg & Tchekhovskoy 2016) and disrupting the accretion flow. Thus, two different regimes, resulting from different Bondi radii, were identified.
@@
Studying a similar problem, Kwan et al. (2023) conducted a series of GRMHD simulations of spherical accretion of magnetized medium with varying initial angular momentum. In contrast to Ressler et al. (2021) and Lalakos et al. (2024), they argued that no persistent jet can form in the zero-angular-momentum case. The inclusion of angular momentum resulted in a powerful jet capable of propagating to large distances. They attributed this to the dependence of the efficiency of the angular momentum transport on the magnetorotational instability.
@@
The diversity and seeming incompatibility of the results in the literature on the properties of jets in low-angular-momentum flows motivates us to reexamine this problem. Spherical accretion of magnetized plasma is also the simplest problem in which to understand magnetic flux accumulation by a central black hole, a key problem in accretion theory given the central role of magnetic fields in angular momentum transport and jet/outflow production. The most idealized version of this scenario involves only a few key parameters: the spin of the BH, the Bondi radius of the surrounding medium, the strength of the magnetic field at the Bondi radius (characterized by plasma-$\beta$) and its orientation relative to the BH spin, and the initial angular momentum (or, equivalently, circularization radius) and its direction. In this study, we restrict our analysis to the case of aligned magnetic, rotation, and BH spin directions but study the effects of different Bondi radii, initial magnetic field strengths, black hole spin, and initial gas angular momentum.
@@
The remainder of the paper is organized as follows. We describe the setup of the GRMHD simulations used in this work in § 2 and present our results in § 3. We first provide an illustrative example of the dependence of the magnetized spherical accretion problem on the chosen Bondi radius in § 3.1. Then, we discuss the dependence of our results on the key parameters of our study in § 3.2; we consider the case of finite initial angular momentum – and the time evolution of the resulting angular momentum – in § 3.3.
@@ 2 Method
We conduct GRMHD simulations utilizing the GPU-accelerated code H-AMR (Liska et al., 2022), which solves the MHD equations in a conservative form in curved space-time. Employing horizon-penetrating Kerr-Schild (KS) coordinates $r_{KS}$, $\theta_{KS}$, $\phi_{KS}$, we concentrate on a nonspinning BH characterized by a dimensionless spin parameter $a=0$, as well as a highly spinning BH with $a=0.95$. In the case of a highly spinning BH, its spin axis aligns with the polar axis. Our results, however, are insensitive to the inclination of the metric relative to the polar singularity of the coordinate system [^1] We demonstrate the insensitivity to the orientation with respect to the pole by tilting the BH spin by $90^{\circ}$ in Appendix A..
We conduct simulations at three resolutions, which we label as small, medium, and large. These correspond to effective grid sizes, $N_{r}\times N_{\theta}\times N_{\phi}$, of $384\times 144\times 256$, $812\times 384\times 512$, and $1536\times 768\times 1024$ cells. We employ “derefinement” near the pole, which efficiently reduces computational costs while maintaining a comparable spatial resolution near both the pole and the equator. The outer boundary, $R_{\rm out}$, is set to $10^{4}r_{g}$ in all simulations unless stated otherwise (the outer boundary of one of the simulations was extended to $10^{5}r_{g}$ to accommodate a Bondi radius of $5000r_{g}$). Our primary focus will be on the medium-resolution runs, as we observe the convergence of the main results with resolution (see Appendix A). Throughout the paper, we measure distance and time in $r_{g}=GM/c^{2}$ and $r_{g}/c$ respectively, where $G$ is the gravitational constant and $M$ is the mass of the central BH.
@@
where $c_{s}$ represents the initial gas sound speed. Initially, a zero-angular-momentum fluid with $\gamma=5/3$ is distributed throughout the simulation domain, characterized by a uniform density $\rho_{0}$ (we set $\rho_{0}=1$ in code units). The internal energy density is also uniform and determined by $r_{b}$:
@@
We set $\beta_{0}=10^{2}$ in the initial conditions of our $r_{b}=100r_{g}$ simulation. In a magnetized Bondi solution with a radial monopole magnetic field, $B^{r}$ undergoes stretching, $B^{r}\propto r^{-2}$. Therefore, to maintain similarity in the plasma-$\beta$ and magnetic flux ($\Phi_{\rm BH}$, described below) near the BH across different runs, we scale it according to the Bondi profile, $\beta\propto r^{3/2}$. Thus, we set $\beta_{0}=10^{3}$ for $r_{b}=500r_{g}$ and $\beta_{0}=3\times 10^{3}$ for $r_{b}=1000r_{g}$. This choice is not critical, however, as our results are in fact insensitive to the specific $\beta_{0}$ value, as long as it is initially above unity. We validate this statement by considering 4 values of $\beta_{0}$ for $r_{b}=500r_{g}$ (see § 3.2): $10$, $10^{2}$, $10^{3}$, and $10^{4}$. This encapsulates our simulations’ three key parameters: spin $a$, Bondi radius $r_{b}$, and initial plasma-$\beta_{0}$. We also discuss the dependence of our results on the initial angular momentum of the gas, parameterized using the circularization radius, $r_{c}$, in § 3.3.
The Bondi accretion rate for an adiabatic index of $\gamma=5/3$ is:
@@ 3 Results
First, we focus on the zero initial angular momentum case and the illustrative comparison of $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$ (with $\beta_{0}=10^{2}$ and $\beta_{0}=10^{3}$ respectively) at $a=0$ and $a=0.95$ in § 3.1. These results are in fact more generic than they might seem because we find a convergence of our results as we increase $r_{b}$ yet further and vary $\beta_{0}$ (§ 3.2). Finally, the impact of angular momentum is discussed in § 3.3.
@@ 3.1 Early evolution
In this section, we discuss the time evolution of our four main cases. We then discuss the role of eruption episodes as well as the magnetic field evolution. The nonspinning case converges with time and reaches a steady state. The highly spinning case initially reaches a similar steady state. However, at a later time, the dynamics of the spinning case with a larger Bondi radius, $r_{b}=500r_{g}$, is governed by the angular momentum generated by the BH feedback; this will be discussed in more detail in § 3.3.
@@ 3.1.1 Temporal evolution
Initially, the accreting plasma undergoes freefall onto the BH, bringing the magnetic flux toward the horizon. Consequently, with the increasing accretion rate at the horizon, plasma properties such as density and velocity conform to the expected Bondi profiles. As long as the magnetic field is dynamically subdominant, $B^{\theta}$ is advected by the radial velocity, while $B^{r}$ experiences stretching according to $B^{r}\propto r^{-2}$. Therefore, the total pressure for a Bondi-like flow profile is given by:
$P_{\rm tot}(r,\theta)=\frac{2\rho_{0}M}{\gamma r_{b}}\left[\left(\frac{r_{b}}{r}\right)^{5/2}+\frac{1}{\beta_{0}}\left[\sin^{2}\theta+\cos^{2}\theta\left(\frac{r_{b}}{r}\right)^{4}\right]\right],$ (9)
where the first and second terms correspond to the thermal and magnetic pressure components respectively. For $\beta_{0}\gg 1$, this yields $P_{B}\ll P_{\rm th}$ close to the equatorial plane $\theta=\pi/2$. At higher altitudes, close to the BH, $P_{B}$ exceeds $P_{\rm th}$ due to its stronger dependence on radius compared to $P_{\rm th}$.
Figure 1: Comparison of the time evolution of the (a) accretion rate $\dot{M}/\dot{M}_{B}$; (b) magnetic flux $\Phi_{\rm BH}$; (c) normalized dimensionless magnetic flux $\phi_{\rm BH}$, and (d) outflow energy efficiency $\eta$ for different $r_{b}$ ($100r_{g}$ and $500r_{g}$ are shown by different colors) and spin values ($a=0$ and $a=0.95$ are represented by thin and thick lines respectively). All quantities are measured at $5r_{g}$. On the right panel ((e)-(h)), a zoom-in of panels ((a)-(d)) is shown. Note the different x-axes for $r_{b}=100r_{g}$ (bottom x-axis) and $r_{b}=500r_{g}$ (top x-axis), represented by the corresponding colors.
@@
At a smaller Bondi radius (green lines), the dimensionless magnetic flux reaches a value indicative of the MAD state, $\phi_{\rm BH}\approx 50$ (c). The energy outflow efficiency, $\eta$, is above $100\%$, demonstrating efficient jet production through the Blandford-Znajek mechanism for $a=0.95$ case (thick green line). For $r_{b}=100r_{g}$ and $a=0.95$, the oscillations in $\Phi_{\rm BH}$, $\phi_{\rm BH}$, and $\eta$ are similar to reconnection-powered eruptions found in MADs that are produced in simulations starting from rotating tori ((f)-(h)).
At a larger Bondi radius of $r_{b}=500r_{g}$, however, the quasi-state dimensionless magnetic flux is smaller than for $r_{b}=100r_{g}$. This trend persists for both nonspinning (thin lines) and spinning (thick lines) cases though the effect is stronger for the spinning BH (thick purple line saturates at a smaller value, compared to the thin purple line in panels (b) and (c)).
@@
At a later time, the spinning BH with a larger Bondi radius (thick purple line) shows an increased $\phi_{\rm BH}$ and efficient energy outflow $\eta$ in the form of a powerful jet. We find such periodic jets to be a consequence of a periodic increase in the angular momentum in the inflow, which will be explored in more detail in § 3.3.
@@
Figure 2: Plasma-$\beta$ in the poloidal plane for runs with different initial parameters $r_{b}$ and $a$; black lines represent magnetic field lines. The first two panels show $a=0$, and the last two show $a=0.95$. (a): $r_{b}=100r_{g},\ \beta=10^{2},\ a=0$; (b): $r_{b}=500r_{g},\ \beta=10^{3},\ a=0$; (c): $r_{b}=100r_{g},\ \beta=10^{2},\ a=0.95$; (d): $r_{b}=500r_{g},\ \beta=10^{3},\ a=0.95$. The snapshots are taken at a time $40000r_{g}/c$ for $r_{b}=100r_{g}$ and $120000r_{g}/c$ for $r_{b}=500r_{g}$.
@@
In Figure 2, we present a side-by-side comparison of plasma-$\beta$ in the same four simulations in the poloidal $r-\theta$ plane with magnetic field lines shown by black lines. The first two columns illustrate the case of a nonspinning BH with $r_{b}=100r_{g}$ (a) and $r_{b}=500r_{g}$ (b). Panels (c) and (d) showcase a highly spinning BH, $a=0.95$, again with $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$, respectively. In both cases where $r_{b}$ is smaller ($100r_{g}$, corresponding to higher gas temperature), a more laminar flow is observed compared to the scenario with $r_{b}=500r_{g}$. Moreover, in the case of a spinning BH, a smaller $r_{b}$ shows the presence of a jet (c), which is not seen in the $r_{b}=500r_{g}$ case (d). The snapshots are taken at the time of $40000r_{g}/c$ and $82000r_{g}/c$ for $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$, respectively.
@@
Figure 3: Three-dimensional rendering of the late-time evolution of plasma-$\beta$ (note that the opacity is decreasing as the plasma-$\beta$ is increasing) and magnetic field lines, plotted on a scale of $250r_{g}$ for $r_{b}=100r_{g}$ (top row, (a)-(b)) and $r_{b}=500r_{g}$ (bottom, (c)-(f)). The origin of magnetic field lines for rendering is fixed at two regions: around the BH and at $\approx 100r_{g}$ from the BH. The corresponding movies can be found for $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$.
@@
In Figure 3, we show a three-dimensional rendering of plasma-$\beta$ (in color) and magnetic field lines for $r_{b}=100r_{g}$ and $r_{b}=500r_{g}$ at $a=0.95$ at a large scale of $500r_{g}$. The magnetic field lines originate from two regions: on a sphere around the BH and approximately $100r_{g}$ from the BH. In all snapshots, the spin axis of the BH points upward. The top row ((a)-(b)) shows two snapshots for the case of $r_{b}=100r_{g}$. Both demonstrate that the powerful jet extends to large distances and is likely kink-unstable beyond the Bondi radius. Magnetic flux tubes, which are remnants of previous eruption episodes we will discuss in detail in the next section, are present around the BH. Panels ((c)-(f)) show four snapshots for the case of $r_{b}=500r_{g}$. At times of $30\times 10^{3}r_{g}/c$ (c) and $100\times 10^{3}r_{g}/c$ (e), the jet is weak and is significantly tilted relative to the vertically oriented spin axis of the BH. Between those times, a jet with $\eta\approx 100\%$ reappeared at a time of $60\times 10^{3}r_{g}/c$ (d). At a time of $160\times 10^{3}$ (f), $\phi_{\rm BH}$ increases again (as seen in Fig. 1), leading to a strong jet with $\eta>100\%$.